The length of a rectangle is 1cm longer than its width. If the diagonal of the rectangle is 4cm, what are the dimensions(the length and width) of the rectangle?
x + (x+1)= 4
Am I on the right track in solving this problem?

Answer by oscargut(2103)

Let L=length and W=width
then L=W+1
and using pitagoras theorem L^2+W^2=4^2
then (W+1)^2+W^2=16
then 2W^2+2W-15=0
so W= (-2+sqrt(124))/4 or W=(-2-sqrt(124))/4 the second solution is negative so
W= (-2+sqrt(124))/4 using L=W+1 you can calculate L